Wavelength at which the spectral radiancy of sun is maximum 500 nm. calculate the temperature of Sun ?
1
Expert's answer
2020-12-14T12:16:14-0500

According to the Wien's displacement law,

"\displaystyle \lambda_{peak}=\frac{b}{T}"

where b ≈ 2898 μm⋅K is Wien's constant.

Therefore,

"\displaystyle T=\frac{b}{\lambda_{peak}} = \frac{2898\cdot 10^{3} \;nm\cdot K }{500 \; nm} = 5796 \;K"


Answer: 5796 K.